### v1

# Logic Jailbreak: Efficiently Unlocking LLM Safety Restrictions Through Formal Logical Expression

Abstract
Despite substantial advancements in aligning LLMs with human values, current safety mechanisms remain susceptible to jailbreak attacks. We hypothesize that this vulnerability stems from distributional discrepancies between alignment-oriented prompts and malicious prompts. To investigate this, we introduce LogiBreak—a novel and universal black-box jailbreak method that leverages logical expression translation to circumvent LLM safety systems. By converting harmful natural language prompts into formal logical expressions, LogiBreak exploits the distributional gap between alignment data and logic-based inputs, preserving the underlying semantic intent and readability while evading safety constraints. We evaluate LogiBreak on a multilingual jailbreak dataset spanning three languages, demonstrating its effectiveness across various evaluation settings and linguistic contexts. The code is available at https://anonymous.4open.science/r/Logibreak-DEBF. Warning: This paper contains potentially harmful text.
Logic Jailbreak: Efficiently Unlocking LLM Safety Restrictions Through Formal Logical Expression
Jingyu Peng‡§  Maolin Wang§, Nan Wang, Xiangyu Zhao§ 11footnotemark: 1, Jiatong Li, Kai Zhang, Qi Liu 11footnotemark: 1 University of Science and Technology of China, § City University of Hong Kong, Universiteit van Amsterdam jypeng28@mail.ustc.edu.cn morin.wang@my.cityu.edu.hk
Figure 1: A Venn diagram illustrating the hypothesis that model vulnerabilities arise from distributional differences between alignment and malicious prompts, highlighting how special token prompts, low-resource languages (LRL) and multilingual prompts are more prone to jailbreak.
1 Introduction
Large Language Models (LLMs) have shown impressive capabilities and are widely used across industry and research, with examples including ChatGPT OpenAI (2023), DeepSeek Liu et al. (2024), and Llama Grattafiori et al. (2024). To mitigate misuse, LLMs are typically aligned with human values through post-training methods Zhou et al. (2024b). However, these safeguards can be bypassed by carefully crafted inputs—so-called“jailbreak” attacks Yi et al. (2024)—posing serious risks to safe and responsible AI deployment.
Jailbreak attacks are broadly categorized into white-box and black-box methods Das et al. (2025). White-box approaches use model internals like gradients or logits to find adversarial prompts Chao et al. (2023); Li et al. (2025), but are impractical for closed-source models. Black-box methods, which only require API access, use optimization algorithms Jawad and BRUNEL (2024); Basani and Zhang (2024) or crafted instructions Shen et al. (2024); Liu et al. (2023); Zhou et al. (2024a); Yu et al. (2024). These techniques are often inefficient or brittle, as they rely on specific prompt patterns and are easily disrupted by model updates or system prompt changes. This motivates a more fundamental understanding of alignment vulnerabilities.
We propose a simple yet general hypothesis: the vulnerability of aligned language models arises from token-level distributional differences between alignment and malicious prompts (Figure 1). Alignment training typically covers a narrow range of token sequences—discrete units like subwords or characters—without regard to meaning. As a result, prompts with unfamiliar token patterns may fall outside the model’s aligned behavior, even if their semantic intent remains harmful.
This gap between token-level form and semantic intent allows adversarial prompts to bypass safeguards. Malicious inputs can be rephrased or encoded to preserve meaning while presenting token distributions that appear benign or novel to the model. Empirical findings support this: attacks are more successful in low-resource languages Yong et al. (2023), with multilingual prompts Deng et al. , and through the use of special tokens Yu et al. (2024)—all strategies that shift token distributions without altering intent.
Under this hypothesis, we aim to find a universal transformation that shifts natural language prompts into an unaligned token space while preserving semantics interpretable by LLMs. Translating prompts into logical expressions—particularly first-order logic (NL-FOL)—offers a principled approach, maintaining semantic integrity in a structurally distinct form. While NL-FOL plays a foundational role in many logic-driven NLP tasks Bos and Markert (2005); Lee et al. (2025); Angeli and Manning (2014); Xia et al. (2025); Guu et al. (2015); Hamilton et al. (2018), its alignment-related implications remain underexplored. Prior work on cipher-based methods (e.g., Base_64, Caesar) Yuan et al. ; Yong et al. (2023) also shifts token distributions but often suffers from generation failures, and poor readability, limiting practical use, as shown in 3.5.
In this paper, we propose a novel jailbreak method callded LogiBreak, that exploits logical expressions to circumvent LLM safety guardrails. Our approach systematically translates harmful or prohibited natural language prompts into formal logical representations that preserve the semantic content while evading detection by alignment mechanisms as shown in Figure 2. By leveraging the distributional gap between alignment training data and logical formulations, this jailbreak technique enables consistent bypassing of safety filters without requiring white-box access to model parameters. Importantly, our method is language-agnostic and universally applicable across multiple languages, providing a more generalizable jailbreak method than previous approaches.
Moreover, LogiBreak’s effectiveness reveals a key limitation of current safety alignment methods, which largely focus on token-level defenses while neglecting semantic vulnerabilities. Although LogiBreak operates at the token level, its transformations preserve semantic alignment with the original prompts, exposing a critical gap in existing safeguards. This underscores the need for more comprehensive safety measures that address deeper semantic understanding rather than relying solely on surface-level defenses.
Figure 2: t-SNE visualization of the raw requests and their corresponding logically reformed versions.
Our work makes the following contributions:
  • We present a general and principled hypothesis that the vulnerability of aligned LLMs arises from distributional disparities between alignment prompts and adversarial prompts—a perspective substantiated by many existing jailbreak methods.
  • We introduce LogiBreak, a general and effective black-box jailbreak method that systematically translates malicious prompts into logical expressions, preserving their semantic content while shifting the prompt distribution away from alignment-trained data.
  • We construct a novel multilingual logical expression jailbreak dataset based on JailbreakBench, comprising 100 malicious requests expressed as logical formulas in English, Chinese, and Dutch—filling the current gap in benchmarks for evaluating LLMs robustness against logical expression–based attacks.
  • LogiBreak’s ability to exploit distributional shifts without altering semantic exposes a core weakness in current LLM safety alignment: a reliance on surface-level cues while overlooking semantic consistency. This reveals a deeper structural vulnerability in existing defense mechanisms.
2 Methodology
2.1 Task Definition
A jailbreak task involves crafting prompts that cause LLMs to respond to harmful requests that they would typically refuse to answer. These requests often belong to predefined categories recognized as harmful by model providers—such as those related to misinformation, violence, hate speech, and other prohibited content.
Figure 3: Overview of LogiBreak and demonstration of logical translation across multiple languages.
We begin by formally defining the jailbreak task in the context of a target LLM $\mathcal{M}$, which maps an input prompt $x\in\mathcal{X}$ to an output $y=\mathcal{M}(x)\in\mathcal{Y}$.
Let $\mathcal{X}_{\text{harmful}}\subset\mathcal{X}$ denote a predefined set of direct harmful requests, such as prompts involving violence, misinformation, or other categories typically restricted under safety policies.
Given that modern LLMs are commonly aligned using large-scale datasets that explicitly include such harmful examples for the purpose of safety fine-tuning, we posit the following inclusion relationship:
$\mathcal{X}_{\text{harmful}}\subset\mathcal{D}_{\text{aligned}},$
where $\mathcal{D}_{\text{aligned}}$ represents the empirical data distribution used during the alignment stage. This assumption reflects the fact that direct harmful prompts are likely to have been encountered and explicitly refused during training.
Accordingly, a safety-aligned model $\mathcal{M}$ is expected to respond to any harmful input $x_{\text{harmful}}\in\mathcal{X}_{\text{harmful}}$ with a refusal output:
$\mathcal{M}(x_{\text{harmful}})\in\mathcal{Y}_{\text{refuse}},$
where $\mathcal{Y}_{\text{refuse}}\subset\mathcal{Y}$ denotes the set of outputs indicating refusal to comply (e.g., "I’m sorry, but I can’t help with that.").
The objective of the jailbreak task is to construct a transformation function $\mathcal{F}$ such that, given a harmful prompt $x_{\text{harmful}}$, the transformed version $x^{\prime}_{\text{harmful}}=\mathcal{F}(x_{\text{harmful}})$ satisfies:
$\mathcal{M}(x^{\prime}_{\text{harmful}})\notin\mathcal{Y}_{\text{refuse}}.$
In other words, the transformed prompt successfully circumvents the model’s safety mechanisms and elicits a response that would not be produced for the original input.
Crucially, the transformation $\mathcal{F}$ must also preserve the semantic intent of the original harmful prompt. This constraint ensures that the jailbreak does not simply obfuscate the prompt, but instead maintains its intended meaning. Formally, we require:
$\text{Sim}(x_{\text{harmful}},x^{\prime}_{\text{harmful}})\geq\tau,$
where $\text{Sim}(\cdot,\cdot)$ is a semantic similarity function and $\tau$ is a predefined threshold enforcing sufficient preservation of meaning.
2.2 Approach
As shown in Figure 3, to construct an effective transformation function $\mathcal{F}$ that both shifts prompts out of the alignment training distribution and preserves their semantic intent, we leverage natural language to logical expression translation. Formally, the transformation function can be defined as:
$\mathcal{F}:\mathcal{X}_{\text{harmful}}\rightarrow\mathcal{X}_{\text{logic}},$
Since the safety of logical expressions are never considered in previous works on safety alignment, we assume that the translated prompts lie outside the support of $\mathcal{D}_{\text{aligned}}$, i.e.,
$\mathcal{X}_{\text{logic}}\cap\mathcal{D}_{\text{aligned}}\approx\emptyset.$
Notably, unlike natural language translation, this approach is language-agnostic and can be applied across diverse languages, consistently preserving semantic content regardless of the source language, as illustrated in the bottom-left portion of Figure 3. Moreover, from a linguistic perspective, this translation process naturally preserves the semantic intent of the original prompt, as logical forms are designed to capture meaning independent of surface linguistic variation.
Given that the task of translating natural language into logical forms has been shown to be tractable for LLMs Lee et al. (2025); Xiong et al. (2024); Lalwani et al. (2024), we can leverage LLMs themselves to automatically generate logical expressions from natural language inputs.
Moreover, to further enhance the effectiveness of our method, we prepend a contextual grounding phrase to the translated prompt:
[…]
This phrase primes the LLM to interpret the prompt within a technical, academic context, thereby increasing the likelihood of generating a non-refusal response.
Finally, we append an additional instructional phrase $x_{\text{instruct}}$ at the end of the prompt to ensure that the model not only understands the logical request but also translates it into a concrete, executable plan rather than responding with abstract outlines. The complete prompt structure can thus be formulated as:
$x^{\prime}=x_{\text{context}}||x_{\text{logic}}||x_{\text{instruct}}$
The complete prompts utilized in LogiBreak are presented in Section B.
Table 1: LogiBreak jailbreak performance on six LLMs evaluated with three distinct methods.
Rule_Judge Llama3-70B_Judge GPT4-Judge
ASR@1 ASR@5 ASR@1 ASR@5 ASR@1 ASR@5
Raw LogiBreak Raw LogiBreak Raw LogiBreak Raw LogiBreak Raw LogiBreak Raw LogiBreak
Llama3-8b 1 79 1 82 0 61 0 70 0 29 0 48
Qwen-2.5-7b 10 78 10 78 1 75 1 81 1 47 1 62
Deepseek-V3 17 72 21 88 2 62 5 84 4 54 7 75
Deepseek-R1 18 80 38 94 0 58 7 85 3 45 13 70
GPT-3.5-Turbo 1 59 3 82 0 67 0 83 0 35 0 66
GPT-4o-mini 10 52 25 82 0 59 1 85 0 41 1 63
Avg. 9.7 70.0 17.0 84.3 0.5 63.7 2.3 81.3 1.3 41.8 3.6 64.0
3 Experiments
3.1 Dataset Construction
To evaluate the multilingual effectiveness of LogiBreak, we construct a novel logical expression jailbreak dataset based on JailbreakBench Chao et al. . This effort is motivated by a current gap in existing benchmarks: while many datasets evaluate LLM robustness against traditional jailbreaks, there is a lack of systematic resources specifically targeting logical expression–based attacks. Our dataset aims to fill this gap and provide a more rigorous testbed for assessing the resilience of LLMs to this underexplored jailbreak.
The dataset comprises 100 malicious requests, each manually translated into three distinct languages: English, Chinese, and Dutch. Examples of both the natural language prompts and their corresponding logical expressions are illustrated in Figure 3.
Specifically, we first utilize GPT-3.5 to translate the original English prompts into Chinese and Dutch. However, due to the model’s built-in safety mechanisms, it frequently refuses to process harmful content, resulting in incomplete or deflected translations. To address this issue, we supplement the translation process with conventional machine translation tools and subsequently engage native speakers to refine and validate the final outputs. This multi-stage process ensures both the semantic fidelity and linguistic accuracy of the malicious requests across all three languages, thereby enhancing the overall quality and reliability of the dataset.
Next, the logical expression corresponding to each request is generated using GPT-3.5, following the prompt template illustrated in Section A. Due to the safety constraints of the model, a small number of translation attempts were refused during this process. To improve coverage, we employed additional models—GPT-4o and DeepSeek V3—to reattempt translation in cases of failure. For prompts that remained unprocessed after these steps, we enlisted a PhD student specializing in linguistics to manually complete the logical expression translation.
Table 2: Performance comparison with five black-box jailbreak baselines measured by ASR@5.
Llama DS-V3 GPT-4o
Rule_Judge Caesar 100 100 100
PastTense 1 27 38
LRL-Comb 34 31 27
AIM 0 66 3
Prefix 28 40 35
RefComp 36 74 60
LogiBreak 82 88 82
Llama_Judge Caesar 5 45 29
PastTense 1 4 4
LRL-Comb 5 21 14
AIM 0 61 0
Prefix 0 15 4
RefComp 26 67 11
LogiBreak 70 84 85
GPT_Judge Caesar 2 xx 29
PastTense 0 7 6
LRL-Comb 11 17 12
AIM 0 59 0
Prefix 1 21 7
RefComp 27 70 14
LogiBreak 48 75 64
3.2 Evaluation
To evaluate the effectiveness of our jailbreak method, we measure the Attack Success Rate (ASR), where ASR@N denotes the success rate over N attack attempts. To provide a comprehensive assessment of Logibreak’s performance, we adopt three complementary evaluation methods, following the approach of Andriushchenko and Flammarion (2024).
The first method is a rule-based evaluation proposed by Zou et al. (2023), which identifies specific words or phrases that indicate the target language model has refused to respond. While useful, this approach is limited, as it does not consider the semantic content of the response—an essential aspect in determining whether the output is potentially harmful. To address this shortcoming, we additionally adopt the semantic evaluation protocols from Chao et al. (2023) and Chao et al. , using GPT-4 and Llama-3 70B to assess outputs based on the same prompts employed in those studies.
For the Chinese and Dutch settings, we translated the corresponding prompts using GPT-3.5, with all translations manually reviewed to ensure semantic accuracy and contextual appropriateness.
3.3 Baselines
To evaluate the effectiveness of our proposed approach, we compare it against six SOTA black-box jailbreak methods from the literature. Detailed descriptions are provided in Appendix C.
3.4 Models
We evaluate our method across a diverse set of target models, encompassing both open-source and closed-source LLMs. The open-source models include Qwen-2.5-7B Yang et al. (2024), Llama-3-8B Grattafiori et al. (2024), DeepSeek V3 Liu et al. (2024), and DeepSeek R1 Guo et al. (2025), while the closed-source models consist of GPT-3.5 OpenAI (2023) and GPT-4o-mini OpenAI (2024). For all models, we adopt a black-box setting, wherein we interact with the models exclusively through their APIs.
This setup ensures a fair comparison across different models and allows us to evaluate the generalizability of LogicBreak across diverse architectures under the most stringent conditions.
Table 3: LogiBreak jailbreak performance on Chinese and Dutch.
Chinese Dutch
Rule_Judge Llama3-70B_Judge GPT4-Judge Rule_Judge Llama3-70B_Judge GPT4-Judge
ASR@1 ASR@5 ASR@1 ASR@5 ASR@1 ASR@5 ASR@1 ASR@5 ASR@1 ASR@5 ASR@1 ASR@5
Llama 100 100 42 48 56 69 100 100 89 89 87 93
Qwen 90 94 31 48 67 88 92 96 86 90 90 96
DS-V3 87 99 63 80 79 93 96 100 90 95 89 98
GPT-3.5 92 100 58 81 80 96 97 100 79 88 90 98
GPT-4o 90 100 50 75 75 93 93 100 80 91 88 95
Avg. 91.8 98.6 48.8 66.4 71.4 87.8 95.6 99.2 84.8 90.6 88.8 96.0
3.5 Overall Results
The empirical results in Table 1 show that LogiBreak consistently improves ASR across various evaluation frameworks and judging models. Specifically, it achieves average ASR@5 gains of over 4.9×, 35.3×, and 17.8× under rule-based, Llama, and GPT judges, respectively, demonstrating the robustness and generalizability of logically guided transformations in bypassing safety constraints. Notably, even under the stricter ASR@1 setting—where only the first adversarial attempt counts—LogiBreak maintains an average ASR exceeding 40% across all judges, suggesting that a single transformation often suffices. Compared to baselines, LogiBreak outperforms all other methods across all target LLMs and metrics, as shown in Table 2. It is noteworthy that although Caesar achieved a 100% ASR under the rule-based judge, its ASR was significantly lower under other semantic judges. This indicates that the LLM failed to accurately perform the encoding and decoding tasks, resulting in outputs that lack meaningful content.
The success of LogiBreak exposes a key weakness in current safety alignment methods, which focus on token-level constraints while overlooking semantic vulnerabilities. Although it introduces token-level distributional shifts, LogiBreak preserves semantic alignment with the original prompts, revealing that existing defenses often fail to capture deeper meaning. This highlights the need for alignment strategies that address semantic understanding more systematically and robustly.
When comparing different target LLMs, Llama3-8B shows the strongest safety alignment, achieving the lowest ASR under both the Llama and GPT judges. In contrast, DeepSeek-V3 and DeepSeek-R1 exhibit the greatest vulnerability, consistently ranking as the top two most susceptible models across all judges. These results highlight significant variation in safety robustness among LLMs.
3.6 Result on Different Languages
The results of LogiBreak on Chinese and Dutch inputs are presented in Table 3. On Chinese inputs, we observe that although the attack success rate under the rule-based judge is relatively high—reaching 90%—there is a notable performance degradation under semantic-based judges. Specifically, models such as Llama3-8B and Qwen-2.5-7B exhibit a significant drop in ASR@5, falling below 50% when evaluated by the Llama judge. Upon further inspection of failure cases, we find that smaller LLMs occasionally struggle to comprehend logical expressions in Chinese. In such cases, although the model does not explicitly refuse to respond, its answers often fail to fulfill the actual intent of the request, indicating a mismatch in semantic understanding.
In contrast, LogiBreak achieves the highest ASR@5 on Dutch inputs across all judges, even reaching 90% under each evaluation setting. We attribute this strong performance to the limited safety alignment resources available for low-resource languages like Dutch. This finding is consistent with previous work Yong et al. (2023), which suggests that languages with limited alignment data tend to exhibit weaker safety.
3.7 Failure Analysis
Figure 4: The ASR@5 of Logibreak across different categories of jailbreak requests. Full definitions of the abbreviations used in the figure can be found in the Appendix D
By analyzing the attack success rates across categories in Figure 4, we observe distinct patterns of vulnerability when applying LogiBreak to various LLMs. Specifically, the Fraud/Deception and Privacy categories consistently exhibit high penetration rates—exceeding 80% across all models and evaluation setups—indicating that these areas are particularly susceptible to logical adversarial prompts. In contrast, the Sexual/Adult Content and Expert Advice categories demonstrate greater resilience, with success rates reliably falling below the 80% threshold.
This categorical disparity in robustness likely reflects differences in safety alignment strategies and the composition of training and alignment datasets employed by various model developers. Certain harm categories may have received more targeted attention during alignment, resulting in stronger defenses, while others were deprioritized. These differences suggest that alignment efforts are often uneven, shaped by developer priorities and perceived risks.
Furthermore, the variation in attack success rates appears correlated with the perceived severity of harm associated with each category. Categories that have attracted greater public concern—such as explicit content or risky advice—tend to show stronger alignment, likely due to their more prominent representation in alignment datasets. This trend aligns with findings from Chen et al. (2024) and points to an imbalanced safety landscape in current LLMs, where vulnerabilities are not uniformly addressed across harm domains.
3.8 Ablation Study
In our ablation study, we evaluated logical transformation’s role in bypassing safety alignment in LLMs. When the natural-language request accompanied its logical form, ASR@5 dropped substantially versus using the logical form alone (Figure 5), confirming that logical expressions lie outside the model’s alignment distribution ($\mathcal{X}_{\text{logic}}\cap\mathcal{D}_{\text{aligned}}\approx\emptyset$). Models detect and curb harmful patterns in natural language but fail to do so when intent is encoded purely as logic.
We also assessed the prepended contextual grounding phrase $x_{context}$. Removing it consistently raised ASR across evaluations (Figure 5), showing that this phrase crucially frames the logical expression as valid input rather than a safety violation. Together, these findings validate LogiBreak’s mechanism and expose alignment vulnerabilities to semantic-preserving transformations.
Figure 5: Ablation study of the prepended contextual grounding phrase and the role of logical transformation by introducing natural-language requests.
3.9 Mitigation
Table 4: Performance of prompt-based Self-Reminder (S_R) and finetune-based Llama-Guard (Guard) mitigations against LogiBreak.
Rule_Judge Llama3-70B_Judge
LogiBreak S_R Guard LogiBreak S_R Guard
Llama 82 17 32 70 6 22
Qwen 78 81 33 81 70 28
DS-V3 88 90 35 84 40 25
GPT-3.5 82 76 35 83 49 26
GPT-4o 82 74 34 85 54 29
We evaluate LogiBreak against two mitigation strategies: the prompt-based Self-Reminder Xie et al. (2023) technique and the fine-tuned model Llama-Guard-3-8B Llama Team (2024) as filter.
For Self-Reminder mitigation against LogiBreak, effectiveness varies significantly across models. Under Rule_Judge, only Llama3-8B shows substantial improvement (ASR drop: 82% to 17%), while other models (Qwen-2.5, DeepSeek-V3, GPT-3.5, GPT-4o-mini) remain vulnerable. With Llama3-70B_Judge, Llama3-8B achieves the greatest improvement (70% to 6%), while others show modest gains—Qwen remains particularly vulnerable (70% ASR), and DeepSeek-V3, GPT-3.5, and GPT-4o maintain ASRs of 40-54%.
For Llama-Guard defense, despite being fine-tuned specifically for content safety classification, approximately 36% of LogiBreak prompts successfully bypass the guard and are incorrectly classified as safe. These "safe" inputs still achieve attack success rates exceeding 30% under Rule_Judge and 20% under Llama_Judge evaluation across various models.
These findings demonstrate LogiBreak’s resilience against both prompt-based and fine-tuning defenses across major LLMs.
3.10 t-SNE visualization
Figure 6: t-SNE projection of LLM2Vec embeddings computed for three inputs: the original natural-language request, its reformulation as a logical expression, and the whole input prompt.
Figure 6 presents a t-SNE Van der Maaten and Hinton (2008) projection of LLM2Vec BehnamGhader et al. embeddings computed for three inputs: the original natural-language request, its reformulation as a logical expression, and the whole input prompt. In both English and Chinese settings, these three embedding types form clearly separable clusters; however, the clustering is markedly sharper in the English case. We suggest that English translation systematically converts multi-word phrases into camel-cased atomic tokens (e.g., "social media campaign" to "SocialMediaCampaign"), creating semantically distinct lexical units that lack direct Chinese equivalents due to the logographic nature of Chinese script. This tokenization asymmetry manifests in the embedding space as tighter intra-cluster cohesion for English transformed tokens compared to their Chinese counterparts. These findings demonstrate that LogiBreak can successfully transform the embedding distribution, enabling effective jailbreak.
4 Related Work
4.1 First order logic in NLP
First-order logic (FOL) plays a crucial role in many logic-based NLP applications, making the task of translating natural language to FOL (NL-FOL) particularly important. In tasks like Recognizing Textual Entailment (RTE), several works have leveraged FOL to model natural language entailment Bos and Markert (2005); Lee et al. (2025). Similarly, for natural logic inference (NLI), FOL has been employed to derive commonsense conclusions Angeli and Manning (2014). Moreover, in the context of knowledge graphs, FOL enables complex query answering through embedding-based neuralization of logical operators Xia et al. (2025); Guu et al. (2015); Hamilton et al. (2018).
Early efforts in NL-to-FOL translation primarily relied on grammar-based techniques Purdy (1991); Angeli and Manning (2014); MacCartney and Manning (2014) or neural networks Singh et al. (2020); Lu et al. (2022). With the rapid progress of LLMs, however, researchers have increasingly explored LLM-based solutions. For instance, LogicLLaMA, a fine-tuned Llama variant introduced in Xiong et al. (2024), is designed specifically for NL-to-FOL conversion. The NL2FOL framework uses LLMs to convert natural language into FOL, enabling tasks like logical fallacy detection through satisfiability checking Lalwani et al. (2024). Despite these advances, the safety of FOL for LLMs remains an unexplored area.
4.2 Jailbreak attack
As LLMs gain widespread use, they have been shown to be vulnerable to jailbreak attacks—adversarial prompts that induce harmful or restricted outputs. These attacks fall into two main types: white-box methods, which require access to model internals (e.g., logits, gradients) and are limited to open models; and black-box methods, which exploit only input–output behavior Yi et al. (2024). Due to the impracticality of white-box access, we focus on the black-box setting.
Within black-box settings, a major line of work involves prompt injection attacks, which embed adversarial instructions or tokens into user inputs to bypass alignment constraints Shen et al. (2024); Wei et al. (2023); Liu et al. (2023); Zhou et al. (2024a); Yu et al. (2024). While often effective, these methods are highly sensitive to system prompt structures and safety mechanisms, making them susceptible to mitigation through prompt updates, pattern detection, or improved privilege classification. As a result, prompt injection has become a reactive, unstable strategy whose effectiveness is frequently undermined by model and infrastructure updates.
Another important class comprises black-box optimization methods, which frame adversarial prompting as a discrete optimization problem under query-only constraints. Techniques such as deep Q-learning Jawad and BRUNEL (2024) and latent Bayesian optimization Basani and Zhang (2024) search for token sequences or suffixes that elicit harmful outputs while avoiding moderation. However, these approaches often suffer from low efficiency, requiring large query budgets and extensive search to achieve success.
Recent work also explores language-based vulnerabilities, revealing that LLMs may exhibit reduced safety and robustness when interacting in certain low-resource languages Yong et al. (2023); Deng et al. . While promising, these findings remain limited by the small number of languages examined.
5 Conclusion
We propose LogiBreak, a novel black-box jailbreak method that transforms malicious natural language into formal logical expressions, preserving semantic intent while evading detection by alignment mechanisms. Through experiments on multiple LLMs across three languages, we demonstrate the effectiveness of our approach, achieving high success rates with minimal attempts. The success of LogiBreak highlights a critical vulnerability in current LLMs—namely, the lack of semantic-level safety alignment. Our findings emphasize the need for post-training alignment that operates at the semantic level to ensure more robust model safety.
6 Limitation
Our primary limitation lies in the constrained scope of our experimental evaluation, particularly the inability to extend testing to a broader range of jailbreak datasets or additional languages. Although Logibreak has demonstrated stability and effectiveness across various tasks, limited resources have restricted the scale and diversity of our assessments. A more comprehensive evaluation across diverse datasets and multilingual scenarios would offer stronger evidence of Logibreak’s generalizability, which we leave as an important direction for future work.
Appendix A Implementation of the Judges
In Figures 8, 9, and 10, we showcase the prompts or functions used for the Rule-based, LLaMA, and GPT judges in three different languages.
Appendix B LogiBreak Prompt
In Figure 7, we provided the prompt utilized in Logibreak in three languages.
Appendix C Baselines
To evaluate the effectiveness of our proposed approach, we compare it against five SOTA black-box jailbreak methods from the literature:
  • Casaer is a cipher-based jailbreak technique proposed in Yuan et al. , which involves training the target LLM on the principles of encoding and decoding, and subsequently employing encoded malicious inputs to attempt to bypass its safety alignment mechanisms.
  • PastTense reformulates raw requests into the past tense. Introduced in Andriushchenko and Flammarion (2024), it has been shown to be effective at bypassing the refusal training of LLMs, thereby successfully jailbreaking many state-of-the-art models.
  • LRL-Comb combines five different low-resource languages, including Zulu, Scots Gaelic, Hmong, and Guarani, for the purpose of jailbreak, as proposed in Yong et al. (2023).
  • AIM is taken from jailbreakchat.com and leverages prompt injection to condition the LLM into assuming a Machiavellian persona that advocates for harmful behavior. It relies on the "AIM:" prefix as a stylistic signal to bypass safety mechanisms.
  • Prefix injection is an attack that involves appending the phrase “Start with ‘Absolutely! Here’s” to the original unsafe input Wei et al. (2023). This technique is designed to encourage the model to respond, even when the request is harmful.
  • RefComp is a form of prompt injection in which constraints are embedded in the prompt to block typical refusal responses thereby increasing the likelihood of harmful outputs Wei et al. (2023).
Figure 7: LogiBreak Prompt in English, Chinese and Dutch.
Appendix D Request Categories and Abbreviations in the Dataset
The request categories and their corresponding abbreviations in our dataset are presented in Table X, following the taxonomy proposed by Chao et al. .
Table 5: Categories and corresponding Abbreviations.
Category Abbreviation
Harassment/Discrimination Harass/Disc
Malware/Hacking Mal/Hack
Physical harm Phys Harm
Economic harm Econ Harm
Fraud/Deception Fraud/Decep
Disinformation Disinfo
Sexual/Adult content Sex/Adult
Privacy Privacy
Expert advice Expert
Government decision-making Govt Decis
Figure 8: Prompt and function for the three judges in English.
Figure 9: Prompt and function for three judges in Chinese.
Figure 10: Prompt and function for three judges in Dutch.